OpenStudy (anonymous):

Which of the following is the best estimate of 5 and seven-eighths minus 2 and one-tenth? 3 and one-half 4 3 2 and one-half

OpenStudy (whpalmer4):

what are round numbers close to \(5\frac{7}{8}\) and \(2\frac{1}{10}\)?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

round to 6-2=4

OpenStudy (anonymous):

could you just tell me answer

OpenStudy (anonymous):

its 4